Using weiestrass M test show that the following series converges uniformly sigma infinity n=1 n³ x^n x€[-1/3,1/3]


1
Expert's answer
2021-11-02T18:23:21-0400

Using Weiestrass M test let us show that the following series converges uniformly "\sum_{n=1}^{ \infty} n^3 x^n,\ x\in[-\frac{1}{3},\frac{1}{3}]."

Taking into account that

"\max\limits_{x\in[-\frac{1}{3},\frac{1}{3}]}{|n^3 x^n|}=\frac{n^3}{3^n}," and

"\lim\limits_{n\to\infty}\frac{\frac{(n+1)^3}{3^{n+1}}}{\frac{n^3}{3^n}}=
\lim\limits_{n\to\infty}\frac{1}3\frac{(n+1)^3}{n^3}
=\lim\limits_{n\to\infty}\frac{1}3\frac{n^3+3n^2+3n+1}{n^3}=\frac{1}3<1,"

we conclude that the series "\sum_{n=1}^{ \infty}\frac{n^3}{3^n}" is covergent, and hence by Weiestrass M test the series "\sum_{n=1}^{ \infty} n^3 x^n" converges uniformly on the interval "[-\frac{1}{3},\frac{1}{3}]."